A year ago, Akashdeep left his job at the Best Bicycle Repair Shop, and opened his own business: the “Even Better Bicycle Repair Shop”. He had been earning $40,000 per year at Best Bicycle. His total revenue at “Even Better” was $95,000. To open his shop, Akashdeep cashed in a $10,000 savings certificate which was earning 2% interest per year. His costs for the year were $25,000 to rent his uncle’s garage and $5,000 for supplies.
a) What were Akashdeep’s total explicit costs for the year?
b) What were Akashdeep’s total implicit costs for the year?
c) What was Akashdeep’s accounting profit for the year?
d) What was Akashdeep’s economic profit for the year?
1
Expert's answer
2020-06-03T11:22:25-0400

Explicit costs are costs out of the pocket that is payment that are actually made.


Implicit cost represent opportunity cost of using resources already owned by the firm, for example working in the business while not getting a formal salary.

a)


"Total Explicit cost=intial cost+rent+supplies"

"10,000+25,000+5,000=40,000\/"


b) "Total implicit cost=earnings forgone +interest foregone"

"40,000+(10,000*2\% )=40,200/"

c)

"Accounting profit=Total revenue-explicit cost"


"95,000-40,000=55,000\/"

d) "Economic profit=Total revenue-explicit cost-implicit cost"

"95,000-40,000-40,200=14,800\/"